OpenStudy (anonymous):

Help please ! Factor the trinomial: 2x2 + 7x + 3 A. (2x - 1)(2x + 3) B. (2x + 3)(x + 1) C. (2x + 1)(x + 3) D. (2x + 1)(2x + 3)

OpenStudy (apoorvk):

C.

OpenStudy (anonymous):

YA U can do it vid spilting middle term C IS CORRECT

OpenStudy (anonymous):

c

OpenStudy (hoblos):

2x2 + 7x + 3 = 2x2 +x + 6x +3 = x(2x+1) + 3(2x+1) = (2x+1)(x+3)